OpenStudy (anonymous):

Calculate the average rate of change for the graphed sequence from n = 2 to n = 4. graphed sequence showing (1, -3), (2, -3.5), (3, -6.75), (4, -10.125), (5, -15.1875), (6, -22.78125)

OpenStudy (anonymous):

-6.625 -03.3125 -3.0 -2.0

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

hi

OpenStudy (nikato):

so look at points (2,-3.5) and (4,-10.125) to find the rate of change, youshould find the slope

OpenStudy (nikato):

sllope formula is \[\frac{ y _{1}-y _{2} }{ x _{2}-x _{2}}\]

OpenStudy (nikato):

plug your numbers in \[\frac{ -10.125-(-3.5) }{ 4-2} =?\]

OpenStudy (anonymous):

-3.3125

OpenStudy (nikato):

yea. thats what i got too